OpenStudy (anonymous):

solve for variable. 3w =4 1/5

OpenStudy (anonymous):

4.2/3=1.4

OpenStudy (anonymous):

i think it is x= 1 2/5

OpenStudy (anonymous):

oh sorry w= that

OpenStudy (anonymous):

-7/5 7/5 63/5 -63/5

OpenStudy (anonymous):

7/5

OpenStudy (anonymous):

7/5

OpenStudy (anonymous):

hah yea someone agrees w/ me :)

OpenStudy (anonymous):

well ta ta I have to go do stuff :) bai

OpenStudy (anonymous):

if you convert 4 1/5 into a fraction you get 21/5 so 3x=21/5 you then divide both side by three or multiply them by 1/3 so \[x=\frac{ 21 }{ 5 }*\frac{ 1 }{ 3 }\] \[x=\frac{ 7 }{ 5 }*\frac{ 1 }{ 1 }\] x=7/5